Your relative wants to pick apples from the tree pictured below. Given that the tree casts a shadow of 30 ft. when your 6 ft. neighbor casts a shadow of 10 ft., how long should a latter be to reach the top of the tree?
Answer by checkley77(12844)

X/30=6/10
10X=180
X=180/10
X=18 FEET IS THE HEIGHT OF THE TREE.